# Local Asymptotic Normality for Mixed Fractional Brownian Motion under High-Frequency Observation

### v2

@@ D.1 Toeplitz trace approximations for $(\widetilde{C}_{n},\widetilde{D}_{n})$
$\widetilde{c}_{n}(\lambda):=\frac{f_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\widetilde{d}_{n}(\lambda):=\frac{\dot{f}_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\lambda\in[-\pi,\pi].$ (78)
@@
$b_{H}(\lambda):=\partial_{H}\log f_{H}(\lambda)=\frac{\dot{f}_{H}(\lambda)}{f_{H}(\lambda)}.$ (79)
@@
$\mathrm{tr}(\widetilde{C}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda+o(n),$ (80)
$\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda+o(n),$ (81)
$\mathrm{tr}(\widetilde{D}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda+o(n).$ (82)
@@
In this $L^{2}$ setting, standard Szegő/Avram-type trace theorems for products of Toeplitz matrices apply to the triangular-array symbols $\widetilde{c}_{n},\widetilde{d}_{n}$ and yield (80)–(82) with $o(n)$ remainder. ∎
@@ D.2 Limits of the trace integrals and the constants $T_{1}(H),T_{2}(H)$
$T_{1}(H):=\frac{1}{2\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)\,d\lambda,\qquad T_{2}(H):=\frac{1}{4\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)^{2}\,d\lambda.$ (83)
@@
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}},$ (84)
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}}\,T_{1}(H),$ (85)
$\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{4\pi}{\sigma^{4}}\,T_{2}(H).$ (86)
@@
Moreover $0\leq\widetilde{c}_{n}\leq\sigma^{-2}$, hence dominated convergence gives (84).
Next, $\widetilde{d}_{n}(\lambda)=\widetilde{c}_{n}(\lambda)\,b_{H}(\lambda)$ and $b_{H}\in L^{1}([-\pi,\pi])$ (log singularity), so $|\widetilde{c}_{n}b_{H}|\leq\sigma^{-2}|b_{H}|$ is integrable and dominated convergence yields (85). Finally, $b_{H}^{2}\in L^{1}([-\pi,\pi])$ and $|\widetilde{d}_{n}|^{2}=\widetilde{c}_{n}^{2}b_{H}^{2}\leq\sigma^{-4}b_{H}^{2}$, hence dominated convergence yields (86). ∎
@@ D.3 Quotien of Two Norms
so $\left\lVert\widetilde{C}_{n}\right\rVert_{F}\asymp\sqrt{n}$. Similarly, (82) and (86) give $\left\lVert\widetilde{D}_{n}\right\rVert_{F}\asymp\sqrt{n}$.
@@ D.4 Joint CLT the same as pure-fGn information matrix
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}}$, (87)
cf. Lemma 1 and (77).
@@
$I_{\mathrm{pure}}(\sigma,H)=\begin{pmatrix}\frac{2}{\sigma^{2}}&\frac{1}{\sigma}\,T_{1}(H)\\[2.84526pt] \frac{1}{\sigma}\,T_{1}(H)&T_{2}(H)\end{pmatrix}$, (88)
with $T_{1}(H),T_{2}(H)$ defined in (83).
@@
$S_{\sigma,n}=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),\qquad R_{H,n}=\frac{\sigma^{2}}{2}\Big(Z^{\top}\widetilde{D}_{n}Z-\mathrm{tr}(\widetilde{D}_{n})\Big),\qquad Z\sim\mathcal{N}(0,I_{n}).$
Using $\mathrm{Var}(Z^{\top}AZ-\mathrm{tr}A)=2\mathrm{tr}(A^{2})$ and Lemmas 2021, we obtain
$\mathrm{Var}\!\Big(\frac{S_{\sigma,n}}{\sqrt{n}}\Big)=\frac{\sigma^{2}}{n}\cdot 2\mathrm{tr}(\widetilde{C}_{n}^{2})=\frac{2\sigma^{2}}{n}\left(\frac{n}{2\pi}\int\widetilde{c}_{n}^{2}+o(n)\right)\to\frac{2\sigma^{2}}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}=\frac{2}{\sigma^{2}},$
@@
$\mathrm{Var}\!\Big(\frac{R_{H,n}}{\sqrt{n}}\Big)=\frac{\sigma^{4}}{4n}\cdot 2\mathrm{tr}(\widetilde{D}_{n}^{2})=\frac{\sigma^{4}}{2n}\left(\frac{n}{2\pi}\int\widetilde{d}_{n}^{2}+o(n)\right)\to\frac{\sigma^{4}}{2\cdot 2\pi}\cdot\frac{4\pi}{\sigma^{4}}=T_{2}(H),$
where the last equality uses (86) and (83).
@@
Using (81) and (85),
$\frac{1}{n}\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})=\frac{1}{2\pi}\int\widetilde{c}_{n}\widetilde{d}_{n}+o(1)\to\frac{1}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}\,T_{1}(H)=\frac{T_{1}(H)}{\sigma^{4}},$
@@
$u_{1}\frac{S_{\sigma,n}}{\sqrt{n}}+u_{2}\frac{R_{H,n}}{\sqrt{n}}=\frac{1}{\sqrt{n}}\Big(Z^{\top}\widetilde{M}_{n,u}Z-\mathrm{tr}(\widetilde{M}_{n,u})\Big),\qquad\widetilde{M}_{n,u}:=u_{1}\sigma\,\widetilde{C}_{n}+u_{2}\frac{\sigma^{2}}{2}\,\widetilde{D}_{n}.$
By Lemma 22, we have $\left\lVert\widetilde{M}_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert\widetilde{M}_{n,u}\right\rVert_{F}\to 0$, so Lemma 2 yields a one-dimensional normal limit for every $u$. Therefore the vector converges jointly to a centered Gaussian with covariance matrix (88). ∎
@@
The limiting covariance (88) coincides with the Fisher information of the pure fGn model in the regime $H<1/2$. The Brownian component enters only through the vanishing regularization $\varepsilon_{n}I_{n}$ in $B_{n}=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}$.

### v3

@@ D.1 Toeplitz trace approximations for $(\widetilde{C}_{n},\widetilde{D}_{n})$
$\widetilde{c}_{n}(\lambda):=\frac{f_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\widetilde{d}_{n}(\lambda):=\frac{\dot{f}_{H}(\lambda)}{\sigma^{2}f_{H}(\lambda)+\varepsilon_{n}},\qquad\lambda\in[-\pi,\pi].$ (D.9)
@@
$b_{H}(\lambda):=\partial_{H}\log f_{H}(\lambda)=\frac{\dot{f}_{H}(\lambda)}{f_{H}(\lambda)}.$ (D.10)
@@
$\mathrm{tr}(\widetilde{C}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda+o(n),$ (D.11)
$\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda+o(n),$ (D.12)
$\mathrm{tr}(\widetilde{D}_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda+o(n).$ (D.13)
@@
In this $L^{2}$ setting, standard Szegő/Avram-type trace theorems for products of Toeplitz matrices apply to the triangular-array symbols $\widetilde{c}_{n},\widetilde{d}_{n}$ and yield (D.11)–(D.13) with $o(n)$ remainder. ∎
@@ D.2 Limits of the trace integrals and the constants $T_{1}(H),T_{2}(H)$
$T_{1}(H):=\frac{1}{2\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)\,d\lambda,\qquad T_{2}(H):=\frac{1}{4\pi}\int_{-\pi}^{\pi}b_{H}(\lambda)^{2}\,d\lambda.$ (D.14)
@@
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}},$ (D.15)
$\int_{-\pi}^{\pi}\widetilde{c}_{n}(\lambda)\widetilde{d}_{n}(\lambda)\,d\lambda$ $\longrightarrow\frac{2\pi}{\sigma^{4}}\,T_{1}(H),$ (D.16)
$\int_{-\pi}^{\pi}\widetilde{d}_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\frac{4\pi}{\sigma^{4}}\,T_{2}(H).$ (D.17)
@@
Moreover $0\leq\widetilde{c}_{n}\leq\sigma^{-2}$, hence dominated convergence gives (D.15).
Next, $\widetilde{d}_{n}(\lambda)=\widetilde{c}_{n}(\lambda)\,b_{H}(\lambda)$ and $b_{H}\in L^{1}([-\pi,\pi])$ (log singularity), so $|\widetilde{c}_{n}b_{H}|\leq\sigma^{-2}|b_{H}|$ is integrable and dominated convergence yields (D.16). Finally, $b_{H}^{2}\in L^{1}([-\pi,\pi])$ and $|\widetilde{d}_{n}|^{2}=\widetilde{c}_{n}^{2}b_{H}^{2}\leq\sigma^{-4}b_{H}^{2}$, hence dominated convergence yields (D.17). ∎
@@ D.3 Quotien of Two Norms
so $\left\lVert\widetilde{C}_{n}\right\rVert_{F}\asymp\sqrt{n}$. Similarly, (D.13) and (D.17) give $\left\lVert\widetilde{D}_{n}\right\rVert_{F}\asymp\sqrt{n}$.
@@ D.4 Joint CLT the same as pure-fGn information matrix
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(1)}\binom{S_{\sigma,n}}{S_{H,n}}=\binom{S_{\sigma,n}}{R_{H,n}}$, (D.18)
cf. Lemma 1 and (D.8).
@@
$I_{\mathrm{pure}}(\sigma,H)=\begin{pmatrix}\frac{2}{\sigma^{2}}&\frac{1}{\sigma}\,T_{1}(H)\\[2.84526pt] \frac{1}{\sigma}\,T_{1}(H)&T_{2}(H)\end{pmatrix}$, (D.19)
with $T_{1}(H),T_{2}(H)$ defined in (D.14).
@@
$S_{\sigma,n}=\sigma\Big(Z_{n}^{\top}\widetilde{C}_{n}Z_{n}-\mathrm{tr}(\widetilde{C}_{n})\Big),\qquad R_{H,n}=\frac{\sigma^{2}}{2}\Big(Z_{n}^{\top}\widetilde{D}_{n}Z_{n}-\mathrm{tr}(\widetilde{D}_{n})\Big),\qquad Z_{n}\sim\mathcal{N}(0,I_{n}).$
Using $\mathrm{Var}(Z_{n}^{\top}AZ_{n}-\mathrm{tr}(A))=2\mathrm{tr}(A^{2})$ and Lemmas 2021, we obtain
$\mathrm{Var}\!\Big(\frac{S_{\sigma,n}}{\sqrt{n}}\Big)=\frac{\sigma^{2}}{n}\cdot 2\mathrm{tr}(\widetilde{C}_{n}^{2})=\frac{2\sigma^{2}}{n}\left(\frac{n}{2\pi}\int^{{\pi}}_{{-\pi}}\widetilde{c}_{n}{(\lambda)^{2}d\lambda}+o(n)\right)\to\frac{2\sigma^{2}}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}=\frac{2}{\sigma^{2}},$
@@
$\mathrm{Var}\!\Big(\frac{R_{H,n}}{\sqrt{n}}\Big)=\frac{\sigma^{4}}{4n}\cdot 2\mathrm{tr}(\widetilde{D}_{n}^{2})=\frac{\sigma^{4}}{2n}\left(\frac{n}{2\pi}\int^{{\pi}}_{{-\pi}}\widetilde{d}_{n}{(\lambda)^{2}d\lambda}+o(n)\right)\to\frac{\sigma^{4}}{2\cdot 2\pi}\cdot\frac{4\pi}{\sigma^{4}}=T_{2}(H),$
where the last equality uses (D.17) and (D.14).
@@
Using (D.12) and (D.16),
$\frac{1}{n}\mathrm{tr}(\widetilde{C}_{n}\widetilde{D}_{n})=\frac{1}{2\pi}\int^{{\pi}}_{{-\pi}}\widetilde{c}_{n}{{({{\lambda}})}}\widetilde{d}_{n}{{({{\lambda}})}}{d\lambda}+o(1)\to\frac{1}{2\pi}\cdot\frac{2\pi}{\sigma^{4}}\,T_{1}(H)=\frac{T_{1}(H)}{\sigma^{4}},$
@@
$u_{1}\frac{S_{\sigma,n}}{\sqrt{n}}+u_{2}\frac{R_{H,n}}{\sqrt{n}}=\frac{1}{\sqrt{n}}\Big(Z_{n}^{\top}\widetilde{M}_{n,u}Z_{n}-\mathrm{tr}(\widetilde{M}_{n,u})\Big),\qquad\widetilde{M}_{n,u}:=u_{1}\sigma\,\widetilde{C}_{n}+u_{2}\frac{\sigma^{2}}{2}\,\widetilde{D}_{n}.$
By Lemma 22, we have $\left\lVert\widetilde{M}_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert\widetilde{M}_{n,u}\right\rVert_{F}\to 0$, so Lemma 2 yields a one-dimensional normal limit for every $u$. Therefore the vector converges jointly to a centered Gaussian with covariance matrix (D.19). ∎
@@
The limiting covariance (D.19) coincides with the Fisher information of the pure fGn model in the regime $H<1/2$. The Brownian component enters only through the vanishing regularization $\varepsilon_{n}I_{n}$ in $B_{n}=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}$.